Direct Labor Cost Budget

Dauber Inc. budgeted production of 27,000 personal journals in 20Y6. Each journal requires assembly. Assume that 20 minutes are required to assemble each journal.

If assembly labor costs $11.5 per hour, determine the direct labor cost budget for 20Y6. Do not round your intermediate calculations but, if required, round your final answer to the nearest dollar.
$
Answer #1

Total minutes = 27,000*20 = 540,000

Converted to hours = 540,000/60 = 9,000 hours

Direct labor cost = 9,000 * 11.5 per hour

= 103,500